# Auto-ARGUE: LLM-Based Report Generation Evaluation

### v3

@@ Abstract
Generation of long-form, citation-backed reports is a primary use case for retrieval augmented generation (RAG) systems. While open-source evaluation tools exist for various RAG tasks, ones tailored to report generation are lacking. Accordingly, we introduce Auto-ARGUE, a robust LLM-based implementation of the recent ARGUE framework for report generation evaluation. We present analysis of Auto-ARGUE on the report generation pilot task from the TREC 2024 NeuCLIR track, showing good system-level correlations with human judgments. We further release a web app for visualization of Auto-ARGUE outputs.
@@ 1 Introduction
As RAG applications have proliferated, interest in automatic evaluation methodologies for RAG tasks has grown in tandem. Many works have positioned themselves as general solutions for RAG evaluation [2, 3, 8, 10], and while some evaluation desiderata are shared across tasks, task-dependent considerations persist.
Here, we focus on report generation (RG), a RAG task that aims to produce a long-form, citation-attributed response to a complex user query. Two key features distinguish RG from related RAG tasks—notably, long-form QA. First, RG strongly foregrounds the identity of the user (or requester): the same query should yield different reports for requesters with different levels of education or domain expertise. Second, the ideal report represents a summary over the entire corpus of the most user-critical information, thus stressing coverage where traditional QA emphasizes mere adequacy of the response.
While the “task-agnostic” RAG evaluation methodologies discussed above could be applied to RG, our work focuses on the ARGUE framework [7]—the only framework designed expressly for RG—and presents three contributions:
@@ 2.1 Framework: ARGUE
Citation Evaluation. Citations are assumed to support only the sentence they are attached to. Sentences may bear $\geq 0$ citations. Relevant citations that attest a sentence are rewarded; non-attesting or missing citations are penalized.
Metrics. ARGUE is agnostic to the magnitudes of its rewards and penalties and to the metrics to be reported based on them. However, Mayfield et al. recommend two metrics—sentence precision and nugget recall—discussed below.
We refer readers to [7] for further details on ARGUE.
@@ 2.2 Implementation: Auto-ARGUE
ARGUE leaves much to implementation, such as the judge, the magnitudes of rewards/penalties, the source of nuggets and relevance assessments, and the metrics to report. Here, we detail the choices we made with Auto-ARGUE.
@@
  • Nuggets. Nuggets may have multiple answers (each of which may be attested by $\geq 1$ document(s)) and come in two varieties: “AND” nuggets, for which all answers must be given, and “OR” nuggets, for which only one answer (of several) is required. Nuggets may also have importance labels (“vital” or “okay”). Answer attestation is assessed per-sentence but answers are aggregated across all sentences to identify correctly answered nuggets.
  • Metrics. Auto-ARGUE implements the two metrics suggested by Mayfield et al. [7]. Sentence precision is the proportion of sentences that are attested by each of their citations. Nugget recall is the proportion of nuggets correctly answered by the report, with a weighted variant that weights nuggets by importance (okay=1.0; vital=2.0). Auto-ARGUE also computes (un)weighted F1 score(s) based on these two metrics, which can serve as an overall score for a report. Auto-ARGUE also outputs several other fine-grained metrics.
@@ 2.3 Visualization: ARGUE-viz
ARGUE-viz is a simple Streamlit [^4] https://streamlit.io app for visualizing Auto-ARGUE outputs for individual runs, including both metrics and fine-grained judgments. Users can toggle between per-topic results and aggregated metrics via radio buttons on a sidebar. Per-topic results display core (e.g. sentence precision, nugget recall) and non-core metrics and statistics (e.g. % relevant citations) for that topic as well as detailed judgment information about the report generated for that topic. Judgment information is displayed via two views: a report view that shows report-level information about supported sentences and correctly answered nuggets and a sentence view that shows similar information at the sentence level (e.g. which nugget answers are (not) attested by that sentence). Collectively, these features enable fine-grained human analysis of errors to facilitate system development.
@@ 3 Case Study: TREC NeuCLIR 2024 Report Generation
We evaluate Auto-ARGUE on the 51 runs from the RG pilot task of the TREC 2024 NeuCLIR track [5], which requires generating English reports from one of three non-English collections (Chinese, Russian, Farsi—17 runs each). Human assessors judged sentence support and nugget recall on reports for the same 21 topics for each run. Each topic has 10-20 nuggets, and assessors also identified documents attesting each answer, thus providing the set of relevant documents.
Separately, we obtain the same metrics from Auto-ARGUE, using Llama-3.3 70B as the LLM judge [4] for D, C, G, and H judgments in Figure 1. B judgments use the human relevance assessments. Since all NeuCLIR nuggets are answerable, E and F judgments are not generated. We obtain system rankings from (a) assessor-based and (b) Auto-ARGUE-based macro-average sentence precision and nugget recall across all topics for each language. We compute agreement between these rankings using (i) Kendall’s tau and (ii) accuracy w.r.t. whether two Wilcoxon tests (one per ranking) for each pair of systems $s_{1}$ and $s_{2}$ agree. Figure 2 presents the results. Broadly, we observe good agreement between the two rankings on both metrics, with particularly strong results on sentence precision. More capable LLM judges could yield still further improvements.
@@ 4 Conclusion
This work has introduced Auto-ARGUE—a robust, configurable, LLM-based implementation of the ARGUE framework for report generation (RG) evaluation—as well as ARGUE-viz—a simple Streamlit application for visualization of Auto-ARGUE outputs. Analysis of Auto-ARGUE on the TREC 2024 NeuCLIR RG pilot task with a relatively small open-source LLM judge (Llama-3.3 70B) reveals good correlations with human judgments on system rankings based on the key metrics of sentence precision and nugget recall. We release both Auto-ARGUE and ARGUE-viz to facilitate future work on RG evaluation.

### v4

@@ Abstract
Generation of long-form, citation-backed reports is a primary use case for retrieval augmented generation (RAG) systems. While open-source evaluation tools exist for various RAG tasks, ones tailored to report generation (RG) are lacking. Accordingly, we introduce Auto-ARGUE, a robust LLM-based implementation of the recently proposed ARGUE framework for RG evaluation. We present analysis of Auto-ARGUE on the RG pilot task from the TREC 2024 NeuCLIR track, showing good system-level correlations with human judgments. We further release a web app for visualization of Auto-ARGUE outputs.
@@ 1 Introduction
As RAG tasks have proliferated, interest in automatic evaluation methodologies for these tasks has grown in tandem. Numerous works have positioned themselves as general solutions for RAG evaluation [2, 3, 8, 10], and while some evaluation desiderata are shared across tasks, many task-dependent considerations persist.
Here, we focus on report generation (RG), a RAG task that aims to produce a long-form, citation-attributed response to a complex user query. At least two important features distinguish RG from related RAG tasks, such as long-form QA. First, RG strongly foregrounds the identity of the user (or requester): the same query should yield different reports for requesters with different levels of education or domain expertise. Second, the ideal report represents a summary over the entire corpus of the most user-critical information, thus stressing coverage where traditional QA emphasizes mere adequacy of the response.
Given these considerations, our work focuses on the ARGUE framework [7]—the only framework designed expressly for RG—and presents three contributions:
@@ 2.1 Framework: ARGUE
Citation Evaluation. Citations are assumed to support only the sentence they are attached to. Sentences may bear $\geq 0$ citations. Relevant citations that attest a sentence are rewarded; non-attesting citations, as well as those missing from sentences judged to require them, are penalized.
Metrics. ARGUE does not mandate the reporting of specific metrics; it merely produces judgments from which metrics can be computed. Mayfield et al. recommend two metrics—sentence precision and nugget recall—discussed below.
@@ 2.2 Implementation: Auto-ARGUE
ARGUE leaves much to implementation, such as the judge, the magnitudes of rewards/penalties, the source of nuggets and relevance labels, and the metrics to report. Here, we detail the choices we made with Auto-ARGUE.
@@
  • Nuggets. Nuggets may have multiple answers (each attested by $\geq 1$ document(s)) and they come in two varieties: AND nuggets, for which all answers must be given, and OR nuggets, for which only one answer (of several) is required. Nuggets may also have importance labels (vital or okay). Answer attestation is assessed per-sentence but answers are aggregated across all sentences to identify correctly answered nuggets.
  • Metrics. Auto-ARGUE implements the two metrics suggested by Mayfield et al. [7]. Sentence precision is the proportion of sentences that are attested by each of their citations. Nugget recall is the proportion of nuggets correctly answered by the report, with a weighted variant that weights nuggets by importance (okay=0.5; vital=1.0). An (un)weighted F1 is also produced based on these two metrics, which can serve as an overall score for a report. Auto-ARGUE further outputs several other fine-grained metrics.
@@ 2.3 Visualization: ARGUE-viz
ARGUE-viz is a simple Streamlit [^4] https://streamlit.io app for visualizing Auto-ARGUE outputs for a run, including judgments and metrics. Users can toggle between run-level and topic-level results via radio buttons on a sidebar. Topic-level results display core metrics (sentence precision, nugget recall, F1) and non-core metrics and statistics (e.g. % relevant citations) as well as detailed information about judgments for the report on that topic. Judgment information is displayed via two views: (1) a report view that shows report-level information about supported sentences and (in)correctly answered nuggets, and (2) a sentence view that shows similar information at the sentence level (e.g. which nugget answers are (not) attested by that sentence). Collectively, these features enable fine-grained human analysis of errors to facilitate system development.
@@ 3 Case Study: TREC NeuCLIR 2024 Report Generation
We evaluate Auto-ARGUE on the 51 runs from the RG pilot task of the TREC 2024 NeuCLIR track [5], which requires generating English reports from one of three non-English collections (Chinese, Russian, Farsi—17 runs each). Human assessors judged sentence support and nugget recall on reports for the same 21 topics for each run. Each topic has 10-20 nuggets, and assessors also identified documents attesting each answer, thus providing (binary) relevance labels.
Separately, we obtain the same metrics from Auto-ARGUE, using Llama-3.3 70B as the LLM judge [4] for D, C, G, and H judgments in Figure 1. B judgments use the human relevance assessments. Since all NeuCLIR nuggets are answerable, E and F judgments are not generated. We obtain system rankings from (a) assessor-based and (b) Auto-ARGUE-based macro-average sentence precision and nugget recall across all topics for each language. We compute agreement between these rankings using (i) Kendall’s tau and (ii) accuracy w.r.t. whether two Wilcoxon tests—one for the assessor-based ranking and one for the LLM-based one—agree on a given pair of runs. Figure 2 presents the results. Broadly, we observe good agreement between the two rankings on both metrics, with particularly strong results on sentence precision. More capable LLM judges could yield even stronger agreement.
@@ 4 Conclusion
This work has introduced Auto-ARGUE—a robust, configurable, LLM-based implementation of the ARGUE framework for report generation (RG) evaluation—as well as ARGUE-viz—a simple Streamlit application for visualization of Auto-ARGUE outputs. Analysis of Auto-ARGUE on the TREC 2024 NeuCLIR RG pilot task with an open-source LLM judge of modest size (Llama-3.3 70B) reveals good correlations with human judgments on system rankings based on the key metrics of sentence precision and nugget recall. We release both Auto-ARGUE and ARGUE-viz to facilitate future work on RG evaluation.